# Multi-Agent Trajectory Prediction with Difficulty-Guided Feature Enhancement Network

### v2

@@ IV-C Analysis and Discussion
Computational Performance. We compared the parameters of each method and the average inference time per scenario on the same device. As shown in Tab.VI, our inference time is slightly higher than LaneGCN and close to the lighter models HiVT-128 and LAformer. However, our method has a much higher prediction accuracy than these two methods. Compared with the methods with the highest prediction accuracy, our method has fewer parameters (note that those marked with * use ensemble strategies), which means that our model achieves high prediction accuracy with lower computational cost. Then, we recorded the inference time on the same experimental machine equipped with an NVIDIA GeForce RTX 3090. The model inference time indicates that the real-time latency of DGFNet is within an acceptable range. Since LAformer uses a two-stage inference approach, its inference time is longer despite having fewer parameters than our method.
V Conclusion
@@
In this paper, we propose a novel Difficulty-Guided Feature Enhancement Network (DGFNet) for muti-agent trajectory prediction. Distinguishing from general future enhancement networks, our model emphasizes filtering future trajectories by masking out and retaining only reliable future trajectories for feature enhancement, which greatly improves the prediction performance. Extensive experiments on the Argoverse 1&2 benchmarks show that our method outperforms most state-of-the-art methods in terms of prediction accuracy and real-time processing. Emulating the predictive habits of human drivers is an intriguing direction for future research. This involves emulating human strategies when encountering different vehicles and making effective assumptions in the presence of incomplete perceptual information.

### v3

@@ III-C Analysis and Discussion
Computational performance. We compared the parameters of each method and the average inference time per scenario on the same device. As shown in Fig.4, our inference time is slightly higher than LaneGCN[21] and close to the lighter models HiVT-128[18] and LAformer[26]. However, our method has a much higher prediction accuracy than these two methods. Compared with the methods with the highest prediction accuracy, our method has fewer parameters, which means that our model achieves high prediction accuracy with lower computational cost. Then, we recorded the inference time on the same experimental machine equipped with an NVIDIA GeForce RTX 3090. The model inference time indicates that the real-time latency of DGFNet is within an acceptable range. Since LAformer uses a two-stage inference approach, its inference time is longer despite having fewer parameters than our method.
IV Conclusion
@@
In this paper, we propose a novel Difficulty-Guided Feature Enhancement Network (DGFNet) for muti-agent trajectory prediction. Distinguishing from general future enhancement networks, our model emphasizes filtering future trajectories by masking out and retaining only reliable future trajectories for feature enhancement, which greatly improves the prediction performance. Extensive experiments on the Argoverse 1&2 benchmarks show that our method outperforms most SOTA methods in terms of prediction accuracy and real-time processing. Emulating the predictive habits of human drivers is an intriguing direction for future research. This involves emulating human strategies when encountering different vehicles and making effective assumptions in the presence of incomplete perceptual information.